Jennifer, Tommy and Linda went to the candy store. Jennifer bought 4 sticks of gum, 10 suckers, and 10 bags of jelly beans. She spent a total of $39.60. Tommy spent $24.90 on 8 sticks of gum, 2 suckers, and 5 bags of jelly beans. Linda bought 9 sticks of gum, 2 suckers, and 8 bags of jelly beans for 33.20. What was the price of each type of candy?
Sticks of Gum: $
Suckers: $
Bags of Jelly Beans: $

Answer by TimothyLamb(4379)

---
x = price per gum
y = price per sucker
z = price per jellies
---
4x + 10y + 10z = 39.60
8x + 2y + 5z = 24.90
9x + 2y + 8z = 33.20
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = price per gum = $1.40
y = price per sucker = $1.10
z = price per bag of jellies = $2.30
---
system matrix
4 10 10
8 2 5
9 2 8
---
inverse of system matrix
-0.032258065 0.32258065 -0.16129032
0.10215054 0.31182796 -0.32258065
0.010752688 -0.44086022 0.38709677
---
determinant of system matrix = -186
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php